OpenStudy (anonymous):

Could anyone show me the steps on how to solve this problem? square root x^2+12x+3=-x

OpenStudy (anonymous):

I got x=-1/4 is this right?

OpenStudy (anonymous):

\[\sqrt{x^2+12x+3}=-x\]

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt {b^2-4ac }}{ 2a }\]

OpenStudy (anonymous):

but its radical expression though o:

OpenStudy (anonymous):

first square both the sides

OpenStudy (anonymous):

I dont understand that part but i know you plug the numbers into the formula. Sorry T: i'm confused.

OpenStudy (anonymous):

squaring both sides will give you x^2+12x+3=x^2 12x=-3 x= -3/12=-1/4